A sum of money is to be distributed amongst A, B, C, and D in the proportion of 5: 2: 1: 4. If D gets Rs. 2000 more than B, then what is the A's share plus C’s share?
1
Expert's answer
2020-04-27T19:14:59-0400

Let the total money be x

then A's share = "\frac5{12}x"


B's share ="\frac2{12}x"


C's share ="\frac1{12}x"


D's share ="\frac4{12}x"


If D gets Rs. 2000 more than B which means

"\frac4{12}x-\frac2{12}x=2000"

"x=12000"


A's share = "12000\times\frac{5}{12}=5000"


C's share ="12000\times\frac{1}{12}=1000"


(A+C)'s share = 5000+1000=6000